Fill 5 odd numbers between 1-15 so that their sum total is 30.
 [ _ + _ + _ + _ + _ = 30]
 

Grade:11

5 Answers

Kartikey Sakare
6 years ago
The answer never can`t be even.If we add 5 odd numbers the anser will be odd only, but the answer here given 30 is wrong.😆#prasad sakare
Ramesh sharma
6 years ago
That`s not possible because 2 odd= 1 even , 5 odd is always odd . So this question is not possible. And if any other answer of this Question exist please tell me
Kartikey Sakare
6 years ago
The answer given 30 is wrong.Becoz sum of 5 odd numbers must be odd so it can`t be answer.for eg: 1+3=4,4+5=9,9+7=16,16+9=25. Therefore odd+odd=evenEven+odd=oddOdd+odd=evenEven+odd=odd
venkatesh
6 years ago
it is not possible because sum of odd number times of odd numbers will be a odd number
so it never be a even number
sahil
6 years ago
it is not possible because sum of any 5 odd no is odd not even like in this case 30 if answer exist reply